OpenStudy (anonymous):

Factor 5a^3 + 15a^2 - 6a - 18

OpenStudy (anonymous):

5a^2(a+3) - 6(a+3) (5a^2-6) (a+3) 5a^2 = 6 a = -3 a^2 = 6/5 a = plus or minus square root of 6/5 AND -3